OpenStudy (red3134):

Given: F = {(0, 1), (2, 4), (4, 6), (6, 8)} and G = {(2, 5), (4, 7), (5, 8), (6, 9), (7, 5)} (F + G) (2) = ?

OpenStudy (danjs):

(F + G)(x) = F(x) + G(x)

OpenStudy (danjs):

(F + G)(2) = F(2) + G(2) = 4 + 5 = 9

OpenStudy (red3134):

@DanJS Ah.... Thank you! I feel kinda dumb now, lol

OpenStudy (danjs):

welcome